I heard that Steve did great work on oplev in Feb 2012.
Here's summary what happened to oplev since then.
Someone changed oplev filters and gain. I couldn't find elog about it. Does anyone know?
Quadrant sum (C1:SUS-XXX_OLSUM) for each optic now and in Feb 2012 is
ITMX ITMY ETMX ETMY BS PRM SRM
2456 14630 1476 14885 3650 4302 2937 counts (now)
1300 14500 900 9000 3500 4000 2600 counts (Feb 6, 2012 elog #6256)
0.025 0.3 0.2 0.2 0.05 0.06 0.04 mW on QPD (Feb 6, 2012 elog #6256)
1350 15000 1500 15500 3500 4000 2600 counts (Feb 23, 2012 elog #6744)
ETMX oplev laser was replaced on May 22, 2012, and quadrant sum was 20500 counts at that time (elog #6656).
Oplev servo openloop transfer functions:
In Feb 2012, gains were adjusted and filter settings are recorded by Steve.
For all pitch OLTF, see elog #6309.
For all yaw OLTF, see elog #6323.
All the filters in Feb is listed in elog #6744.
Filters now are messed up, as Jamie pointed out in elog #6743.
Below is the current filter settings.
I turned ELP and RLP filters on, which wasn't on to cut-off noises at higher frequencies.
I left resonant gains of ETMs because I don't know what they are for.
I put ELP35 for ITMs, BS, PRM and SRM. I put RLP80 for BS, PRM and SRM.
I will leave ELP35 off for BS and SRM because they oscillate currently. ELP50 and ELP40 is on for a substitution. I will readjust them soon.
I don't know who changed all gains (except for PRM, which I adjusted in elog #6952). It doen't look like it is because of change in quadrant sum.
I also don't know who deleted 3.3 Hz resonant gain for BS.
I put all similar filters in same place to make it organized. Now, basic fitlers are organized. We may need some resonant gains for each optics.